OpenStudy (anonymous):

Given the linear equation , find the y-coordinates of the points (-9, ), (-3, ), and (6, ).

OpenStudy (anonymous):

wich equation

OpenStudy (anonymous):

y=-2/3x-2

OpenStudy (id21):

x=-9 y=-2*(-9)/3-2=6-2=4 x=-3 y=-2*(-3)/3-2=2-2=0 x=6 y=-2*6/3-2=-4-2=-6